

********************************************************************
UARS WEEKLY STATUS REPORT - 23 APR 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
NOTHING RECEIVED.
***********
PEM
***********
NOTHING RECEIVED.
***********
HALOE
***********
HALOE continues successful operations with all engineering parameters
within nominal operating range. Science data results are of excellent quality.
***********
MLS
***********
NOTHING RECEIVED.
***********
HRDI
***********
NOTHING RECEIVED.
***********
WINDII
***********
WINDII performed normally during its operating period of
April 13 - 19. Its next operation is scheduled for April
25 - 29. Data reprocessing is proceeding well, and is
up to day 359.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF Y2K end to end testing is scheduled for Thursday April 29, 1999.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
1 day HRD (4.4) PROD (L1-L3) (1R-1C=0L)
50 days PEM (3.32) PROD (OA-L2) (760R-528C-1F=231L)
131 days WIN (5.11) PROD (LOG) (2014R-1192C=822L)
71 days WIN (5.11) PROD (L1-L3) (2014R-263C=1751L)
The CDHF plans to start testing CLAES V7.0 formal software delivery.
SOLSTICE Wall Clock time increase. Analyze SOLSTICE debug step 2 test
production job for "general unrecoverable error".
***********
FOT
***********
REPORTING PERIOD
This report covers 10 April 1999 (Orbit 41410, GMT Day 100) through 16 April
1999 (Orbit 41513, GMT Day 106).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
MLS 100/2314 41424 Off
ACRIM 101/0200 41426 On
SUSIM 101/1257 41433 On
PEM 102/2242 41454 AXIS-1 Off
WINDII 103/0208 41456 On
PEM 107/0040 41515 AXIS On
The beta angle ranged between the angles of 41.4 degrees to 46.5 degrees
for this report period. The beta angle is now increasing toward a maximum
angle of 2.9 degrees on 30 April 1999 (DOY 120).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.2 to 1.6 degrees C. Additional battery performance data
for each battery on 09 April 1999 and 16 April 1999 are provided in
Appendix D.
The clock error ranged from -10.9 to 11.7 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 01
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 05
% ACRIM 01
% MLS 01
% PEM 03
% SUSIM 01
% WINDII 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 04
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/054 - 2287.496923
Transponder B: 99/055 - 2287.498750
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/105 show the observatory data loss to be
22 hours, 47 minutes, 46.804 seconds (an increase of 41.102 seconds since
last report period). This is a 0.0345 percent data loss which equals a
99.9655 percent data capture for the mission.
The increase of 41.102 seconds occurred over four (4) days:
7 0.576 seconds on Day 100 due to data hit,
7 30.112 seconds on Day 102 due to degraded data (reason unknown),
7 1.408 seconds on Day 103 due to data hits, and
7 9.024 seconds on Day 105 due to insufficient overlap.
The FOT is still awaiting the delivery of the 8-week Ground Track Y2K test
file from FDF. This file was expected to be delivered this week, but was
not received as of the end of this reporting period.
A new version of RTS 31 was built on 15 April. The purpose of the new RTS
was to further refine the PMON controlled operations of the PEM instrument.
From the data observed by the FOT after the uplink of this latest RTS, it
appears that the PEM is finally operating as desired via PMON.
The FOT was required to delete and re-add all of their confirmed TDRS
events spanning days 106 through 115 due to an NCC database problem. This
was a lengthy task, as all of the new events had to be built to reflect the
exact AOSs and LOSs of the original events.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT and the PEM team have defined an RTS that successfully operates PEM
EP6 under PMON control. On Wednesday, 14 April, PDB-10 was disabled and
its three command added to the beginning of RTS-31 and the EP6 responded
correctly.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT - No change
UARS PROPULSION MODULE B/D THRUSTER LINE FILL - No change
MEETINGS
FOT supported the weekly IMOC meeting on 14 April 1999. Tim Christy spoke
about the Blossom Point project that he used to work on. It was a Navy
Research Lab multi-satellite operation, so it gave us some more information
on how to conceptualize the IMOC.
FOT attended the monthly battery teleconference. The next battery
teleconference is scheduled for Tuesday, 11 May 1999 at 1:00 PM EDT.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver (reverse-to-forward) is scheduled for 30
April 1999 (DOY 120).
PERSONNEL CHANGES
John Speer has announced his intent to retire effective 30 April 1999.
Pete Thomas, Offline Engineer, announced his resignation effective 21 April
1999 to pursue other opportunities.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 16 April 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
41417 100/1204 99-130 OBC Clk Adj fr 0.00 to 20.6 msec/day
41421 100/1732 99-131 PMON Enabled Comp Htr 18 A&B sides
41424 100/2314 99-132 MLS Power OFF
41426 101/0200 99-133 ACRIM Power ON
41433 101/1257 99-134 SUSIM Power ON
41436 101/1806 99-131 PMON Disable Comp Htr 18 B side
41454 102/2242 99-135 PEM Power OFF AXIS1
41454 102/2242 99-136 OBC Clk Adj fr 20.6 to 0.00 msec/day
41456 103/0208 99-137 WINDII Power ON
41482 104/1956 99-138 OBC Clk Adj fr 0.00 to 20.6 msec/day
41483 104/2135 99-139 PEM RTSPEM Test
41515 107/0040 99-107 PEM AXISPower ON
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 16 April 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
SORLBXTMPR Y-LO Logic_Box Temperature continuous
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 16 April 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
41417 100/12:11 (10 Apr 99) from 00.0 to 20.6 msec/day
41454 102/22:50 (12 Apr 99) from 20.6 to 00.0 msec/day
41482 104/20:05 (14 Apr 99) from 00.0 to 20.6 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 16 April 1999
09 Apr 1999 - GMT Day 099 Beta = 41.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -44.8/-67.2 -3.05/-4.31 +0.0/ +0.0 16.6 0.000 0.0
2 +100.8/0.0 +6.11/+5.15 +28.8/-15.2 24.5 1.022 21.9
3 +22.4/-50.4 +5.15/+3.87 +28.4/-16.0 24.5 1.025 21.8
16 Apr 1999 - GMT Day 092 Beta = 46.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -56.0/-78.4 -3.05/-4.31 +0.0/ +0.0 16.6 0.000 0.0
2 +78.4/-5.6 +6.44/+4.51 +27.2/-16.0 24.3 1.019 22.1
3 +33.6/-33.6 +5.15/+3.87 +27.2/-16.8 24.2 1.023 21.8
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 16 April 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-017 FOT 41421 100/1732 N/A File Transfer PENDING
99-016 FOT 41437 101/1915 14170 Generic Late Acquisition
99-016 FOT 41473 104/0532 14170 Generic Late Acquisition
99-018 FOT 41496 105/1918 20759 Tlm hit I & Q
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-017 FOT 41421 100/1732 N/A File Transfer PENDING
99-016 FOT 41437 101/1915 14170 Generic Late Acquisition
99-016 FOT 41473 104/0532 14170 Generic Late Acquisition
99-018 FOT 41496 105/1918 20759 Tlm hit I & Q
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-016 FOT see list Apr 99 14170 Generic Late Acq
SEVNTFW99/16